In concentration of aqueous solutions find mass fraction and mole fraction using volume fraction of 0.8629 specific gravity of .4040 and MW solute of 59.00
1
Expert's answer
2021-10-05T01:44:10-0400

0.8629×0.4040= 0.3486


0.3486/59= 0.006g

=6/1000g